OpenStudy (anonymous):

Solve for the indicated letter. f=6g for g

OpenStudy (anonymous):

f/6 = g

OpenStudy (anonymous):

Rachel....you know how to divide, right?

OpenStudy (anonymous):

of course

OpenStudy (saifoo.khan):

simply cross multiply.

OpenStudy (anonymous):

saifoo, lol...that's 6 not g

OpenStudy (anonymous):

so confused. LOL

OpenStudy (anonymous):

so that is the answer g=f/g ?

OpenStudy (saifoo.khan):

lol, My bad! g = f/6

OpenStudy (saifoo.khan):

that g became 6

OpenStudy (anonymous):

lol ok...so g=f/6 ?

OpenStudy (anonymous):

Yes, that is correct

OpenStudy (saifoo.khan):

Yeah!

OpenStudy (anonymous):

SWeet! Thanks! Both of you!

OpenStudy (anonymous):

wait so is that f over 6 like a fraction?

OpenStudy (saifoo.khan):

g = f -- 6

OpenStudy (anonymous):

ok so like a fraction! GOT IT! lol

OpenStudy (saifoo.khan):

Yeah!